# A Survey of 3D Reconstruction with Event Cameras: From Event-based Geometry to Neural 3D Rendering

### v1

@@ 7 Research Gaps & Future Direction
Non-frame-transferring synthetic event datasets: Currently, the main modelling platforms do not support event-based 3D modelling. While simulation tools such as ESIM [76] and Video-to-Event [24] exist, the event streams they generate rely on brightness changes between consecutive image frames, which introduces noticeable frame-based artefacts and fails to capture the sparse and asynchronous nature of real event data. In the future, simulating realistic event cameras within 3D modelling environments — enabling the collection of highly realistic event streams with perfect ground truth — holds great potential for deep learning across all event-based tasks.
@@
Reconstruction of low-texture and non-lambertian objects: Reconstructing 3D scenes with low-texture or non-Lambertian surfaces (e.g., glass or mirrors) remains highly challenging and is rarely studied. While the event camera’s sensitivity to photometric changes could potentially complement traditional cameras in such tasks, it may also be adversely affected. To date, only a few studies have explored using event cameras for photometric stereo of these types of objects [107]. Future research should investigate the effectiveness of event cameras in this context and, where applicable, design specialised algorithms tailored to reconstruct these challenging materials.
Hardware and Synchronisation Constraints: The performance of event-driven pipelines is often bounded by hardware limitations, including the event camera’s resolution, timestamp precision, and bandwidth. In multimodal systems, accurate synchronisation between events, frames, IMU, and structured light sources is difficult to achieve and critical for fusion-based reconstruction. Even small pose errors can propagate and degrade 3D structure estimation [29].
@@
Underexplored Modalities: While structured light and D-RGB sensors have been successfully combined with event cameras, modalities like LiDAR, polarisation, or event-based ToF remain underexplored. Their integration could provide additional priors to improve reconstruction under adverse conditions. Similarly, spatio-temporal fusion techniques offer promising opportunities but require careful synchronisation and encoding strategies [20].
Exploring Broader Downstream Applications: Due to the limitations of existing datasets, current research on event-based 3D reconstruction remains constrained to scenarios supported by available benchmarks. However, there are many foreseeable downstream applications yet to be fully explored. For instance, mounting event cameras on unmanned aerial vehicles (UAVs) could enable the reconstruction of large-scale objects or architectural structures under challenging conditions such as high-speed motion or dynamic lighting. Additionally, event-based 3D reconstruction holds great promise in robotics for fast, low-latency perception and navigation in cluttered or fast-changing environments. In autonomous driving, event cameras can provide reliable geometry in high dynamic range scenes or during rapid vehicle motion, complementing LiDAR and conventional cameras. In augmented and virtual reality, event-driven depth sensing may enable energy-efficient and low-latency 3D interaction. Furthermore, industrial applications such as quality inspection of reflective or high-speed moving parts, as well as medical imaging under low-light or non-invasive conditions, represent other promising directions. Future work should focus on expanding the scope of reconstruction tasks and datasets to facilitate progress in these real-world scenarios.
8 Conclusion
@@
This survey presents a comprehensive and systematic review of 3D reconstruction techniques based on event cameras, which are emerging as a powerful alternative to conventional vision sensors in challenging environments. We categorised the literature by input modality — including stereo, monocular, and multimodal systems — and further by reconstruction strategy, ranging from geometry-based and deep learning-based pipelines to recent advances in neural rendering using Neural Radiance Fields and 3D Gaussian Splatting. Methods with a similar research focus were organised chronologically into the most subdivided groups. Through detailed comparisons and timeline visualisations, we revealed the evolution and diversification of event-driven 3D reconstruction. We also compiled a list of publicly available datasets to support reproducibility and benchmarking. Despite the progress, we identified critical research gaps in dataset standardisation, event representation design, dynamic scene modelling, real-time deployment, etc. As event cameras continue to mature, we anticipate further breakthroughs in both theoretical modelling and practical applications, particularly under extreme motion and illumination conditions. We hope this survey provides a useful foundation for new researchers and a roadmap for advancing the field of event-based 3D reconstruction.

### v2

@@ VII Research Gaps & Future Direction
Non-frame-transferring synthetic event datasets: Currently, the main modelling platforms do not support event-based 3D modelling. While simulation tools such as ESIM [105] and Video-to-Event [191] exist, the event streams they generate rely on brightness changes between consecutive image frames, which introduces noticeable frame-based artefacts and fails to capture the sparse and asynchronous nature of real event data. The latest simulator for event cameras is showing the possibilities of creating such datasets [192, 193]. In the future, simulating event cameras within 3D modelling environments, which enables the collection of highly realistic event streams with perfect ground truth, holds great potential for deep learning in 3D reconstruction tasks.
@@
Reconstruction of object with challenging materials: Reconstructing 3D scenes with low-texture or non-Lambertian surfaces [196, 197] (e.g., glass or mirrors) remains highly challenging and is rarely studied. While the event camera’s sensitivity to photometric changes could potentially complement traditional cameras in such tasks, it may also be adversely affected. To date, only a few studies have explored using event cameras for photometric stereo of these types of objects [198]. In addition, dynamic and non-rigid objects introduce further complexity [199], as their continuously changing geometry violates the assumptions of many static reconstruction methods. Recent works in event-based dynamic NeRF [125, 153] show potential in this direction, but general solutions remain underexplored. Future research should investigate the effectiveness of event cameras in this context and, where applicable, design specialised algorithms tailored to reconstruct these challenging materials.
Hardware and synchronisation constraints: The performance of event-driven pipelines is often bounded by hardware limitations, including the event camera’s resolution, timestamp precision, and bandwidth. In multimodal systems, accurate synchronisation between events, frames, IMU, and structured light sources is difficult to achieve and critical for fusion-based reconstruction [200]. Even small pose errors can propagate and degrade 3D structure estimation [71].
@@
Underexplored modalities: While structured light and RGB-D sensors have been successfully combined with event cameras [7], modalities like LiDAR, polarisation, or event-based ToF remain underexplored. Their integration could provide additional priors to improve reconstruction under adverse conditions. Similarly, spatio-temporal fusion techniques offer promising opportunities but require careful synchronisation and encoding strategies [149].
Exploring broader downstream applications: Due to the limitations of existing datasets, current research on event-based 3D reconstruction remains constrained to scenarios supported by available benchmarks. However, there are many foreseeable downstream applications yet to be fully explored:
  • UAVs: Mounting event cameras on unmanned aerial vehicles (UAVs) could enable the reconstruction of large-scale objects or architectural structures under challenging conditions such as high-speed motion or dynamic lighting.
  • Robotics: Event-based 3D reconstruction holds great promise in robotics for fast, low-latency perception and navigation in cluttered or fast-changing environments.
  • Autonomous Driving: Event cameras enable stable 3D perception in scenarios where conventional cameras struggle, such as nighttime, tunnels, or strong backlighting. In autonomous driving, they offer reliable geometric sensing under high dynamic range conditions or rapid motion, serving as a robust complement to LiDAR and standard cameras.
  • VR: In augmented and virtual reality, event-driven depth sensing may enable energy-efficient and low-latency 3D interaction.
  • Cultural Heritage Scanning: Event cameras may enable non-invasive 3D reconstruction of fragile artefacts and heritage sites under low-light or vibration-sensitive conditions, offering a safe alternative for digitising and monitoring valuable cultural assets.
  • Other industrial applications, such as quality inspection of reflective or high-speed moving parts, as well as medical imaging under low-light or non-invasive conditions, represent other promising directions.
VIII Conclusion
@@
This survey presents a comprehensive and systematic review of 3D reconstruction techniques based on event cameras, which are emerging as a powerful alternative to conventional vision sensors in challenging environments. We categorised the literature by input modality, including stereo, monocular, and multimodal systems, and further by reconstruction strategy, ranging from geometry-based and deep learning-based pipelines to recent advances in neural rendering using Neural Radiance Fields and 3D Gaussian Splatting. Methods with a similar research focus were organised chronologically into the most subdivided groups. Through detailed comparisons and timeline visualisations, we revealed the evolution and diversification of event-driven 3D reconstruction. We also compiled a list of publicly available datasets to support reproducibility and benchmarking. Despite the progress, we identified critical research gaps in dataset standardisation, event representation design, dynamic scene modelling, real-time deployment, etc. As event cameras continue to mature, we anticipate further breakthroughs in both theoretical modelling and practical applications, particularly under extreme motion and illumination conditions. We hope this survey provides a useful foundation for new researchers and a roadmap for advancing the field of event-based 3D reconstruction.
Author contributions
Chuanzhi Xu proposed the review topic, collected most of the referenced papers, and conducted and structured the majority of the work. He also created the majority of the tables and figures. Haoxian Zhou contributed significantly to the writing of several sections and produced multiple figures. Langyi Chen wrote Section 6 on datasets. Haodong Chen and Qiang Qu carefully reviewed the manuscript, and Qiang Qu additionally contributed by revising the paper, coordinating the project, and managing subsequent communications. Ying Zhou, Vera Chung, and Weidong Cai supervised the project, provided necessary materials, and offered general guidance and feedback on the manuscript.
Declaration of competing interest
The authors have no competing interests to declare that are relevant to the content of this article.